OpenStudy (anonymous):

if you have a triangle ABC. AB=6 cm, BC=7 cm and AC=x. the angle CAB=60 degrees. how long is x? Ps. you are not allowed to use sine rule only cosine rule.

OpenStudy (campbell_st):

draw the picture  then by the cosine rule \[7^2 = 6^2 + x^2 - 2 \times 6 \times x \times \cos(60)\] which likes like you'll have a quadratic to solve.